# Supervised Graph Contrastive Learning for Gene Regulatory Network

### v3

@@ 5 Experiments
Figure 2: Overview of Downstream Tasks Used for Benchmark of GRNs Across Three Cancer Types: Breast, Lung, and Colorectal Cancers. The node-level tasks involve classifying genes into functional categories such as Biological Process [BP.], Cellular Component [CC.], and cancer relevance [Rel.]. The graph-level tasks include patient-level survival risk prediction [Hazard] and disease subtype classification [Subtype], the latter being specific to breast cancer. Mean pooling is applied to obtain graph-level representations.
Table 1: Description of downstream task
@@
[Subtype]: Disease subtype prediction Classification (5 groups) Subtype accuracy
In this study, we formulated SupGCL by naturally extending the loss function of conventional GCL, based on a contrastive learning framework using KL divergence. Furthermore, we clarified the relationship between SupGCL and conventional GCL through the temperature parameter. This chapter verifies the effectiveness of the proposed method using actual gene regulatory networks (GRNs) from cancer patients and augmented GRNs based on gene knockdown experiments.
@@ 5.1 Benchmark of Gene Regulatory Networks
We evaluated the proposed method through the following procedure. First, based on gene expression data from cancer patients, we constructed patient-specific GRNs. Similarly, we constructed teacher GRNs using gene knockdown experiment data. Then, pre-training was performed on the proposed method using both the patient-specific and teacher GRNs. Subsequently, the performance of the downstream tasks, such as classification accuracy and regression performance, was evaluated using the pre-trained models and compared against comparative methods. Finally, we visualized the latent representations at both the node and graph levels extracted from the trained models.
@@
: Directly performs classification or regression for downstream tasks without any pre-training.
@@
: [19]: Graph representation learning method based solely on graph reconstruction.
@@
: [13]: Graph contrastive learning using positive pairs between graphs.
@@
: [15]: Node-level graph contrastive learning method.
@@
: [17]: Node-level GCL that leverages representation scattering in the embedding space.
Table 2: Finetuning result of node-level downstream task.
@@
Table 3: Finetuning result of graph-level downstream task.
@@
To evaluate the performance of SupGCL, we conducted benchmark evaluations using real-world datasets. For constructing patient-specific GRNs, we used cancer cell sample data from The Cancer Genome Atlas (TCGA). For constructing teacher GRNs, we used gene knockdown experiment data from cancer cell lines in the Library of Integrated Network-based Cellular Signatures (LINCS). The TCGA dataset  [20] and the LINCS dataset  [21] are both large-scale and widely-used public platforms providing gene expression data from cancer patients and cell lines, respectively.
We used normalized count data from the TCGA TARGET GTEx study  [22] provided by UCSC Xena. For LINCS, we used normalized gene expression data from the LINCS L1000 GEO dataset (GSE92742)  [23].
@@
The TCGA dataset also includes survival status and disease subtype labels associated with each gene expression profile. Additionally, each gene was annotated with multi-labels based on Gene Ontology  [24] — Biological Process (metabolism, signaling, cell organization; 3 classes), and Cellular Component (nucleus, mitochondria, ER, membrane; 4 classes). We also used the OncoKB  [25] cancer-related gene list to assign binary relevance labels. These labels were used for downstream tasks. Details are provided in Appendix C.
@@
To estimate the network structure of each GRN from gene expression data, we used a Bayesian network structure learning algorithm based on nonparametric regression with Gaussian noise  [26]. For each experiment, gene expression values were used as node features, while edge features were defined as the product of estimated regression coefficients and the parent node’s gene expression  [27]. This structure estimation was performed per cancer type per dataset using the above algorithm. Further details are provided in Appendix D.
Figure 3: t-SNE visualization of pre-trained embeddings on breast cancer GRNs. The top row shows the node-level embedding space for individual genes, and the bottom row shows the corresponding graph-level readout features for each patient’s network.
@@ 5.2 Result 1: Evaluation by Downstream Task
Graph-level tasks (hazard prediction, subtype classification) used survival and subtype labels from TCGA. Note that subtype classification was conducted only for breast cancer. Node-level tasks (Biological Process - BP, Cellular Component - CC, and cancer relevance - Rel.) used gene-level annotations from Gene Ontology and OncoKB.
Details of these downstream tasks are summarized in Table 1.
@@
For all methods including the SupGCL and conventional methods, we used the same 5-layer Graph Transformer architecture  [28]. Hyperparameters for pre-training were tuned with Optuna  [29], and the model was optimized using the AdamW optimizer  [30]. All training runs were performed on a single NVIDIA H100 SXM5 GPU. Additional experimental details can be found in Appendix E.
@@
For node-level tasks, many existing methods did not show much improvement over without-pretrain, whereas SupGCL consistently demonstrated strong performance. This suggests that SupGCL effectively captures biologically meaningful GRN representations suitable for these tasks. In graph-level tasks like hazard and subtype prediction, while some existing methods showed marginal improvement over without-pretrain, SupGCL achieved significantly higher performance.
5.3 Result 2: Latent Space Analysis
Using the pre-trained models, we visualized the embedding spaces derived from the breast cancer dataset (see Figure 3). Other similar visualizations are shown in Appendix F.
The top row of Figure 3 shows the node-level embeddings. Colors represent one of the three single-label Biological Process annotations (metabolism, signaling, or cell organization). Across all models, no clear clustering was observed based on labels. However, GraphCL’s embeddings were notably different from others, with signs of latent space collapse. Detailed observations are provided in Appendix F. This suggests that GraphCL may be more suited to graph-level discrimination than node-level representation.
Compared to GraphCL, the other models showed more dispersed embeddings, though without distinct clustering.
The bottom row of Figure 3 illustrates the graph-level embeddings, colored by disease subtype. From this, it is evident that GAE and GraphCL fail to separate subtypes in the embedding space. In contrast, GRACE and SGRL showed moderate separation between the Basal subtype and others. SupGCL displayed the clearest separation, indicating its ability to better learn subtype-specific network representations.
@@ 6 Conclusion
A limitation of this paper is that the effectiveness of the proposed method has only been confirmed in situations where the teacher GRNs were constructed from knockdown experiments of the same cancer type.
As future work, we plan to expand the target cancer types and develop a large-scale, general-purpose SupGCL model that can operate across multiple cancer types.
Acknowledgments and Disclosure of Funding
This work was supported by JST Moonshot R&D Grant Number JPMJMS2021, JPMJMS2024, and JSPS KAKENHI Grant Number 25K00148. This work used computational resources of the supercomputer Fugaku provided by RIKEN through the HPCI System Research Project (Project IDs: hp150272, ra000018).
@@ Appendix A Detail Proof of Theorem 1
Since the loss function of contrastive learning (1) is fomulated as the KL divergence between the joint distributions $p(x,y)$ and $q(x,y)$, under the assumption that $p(x)$ and $q(x)$ follow uniform distributions. Then, the following derived for $\mathrm{Loss}_{\rm I-CON}$:
@@
$=\mathrm{E}_{(a,b)\sim p_{\phi}(a,b)}\Big{[}D_{\mathrm{KL}}(p_{\phi}(i,j|a,b)|q_{\phi}(i,j|a,b))\Big{]}+D_{\mathrm{KL}}(p_{\phi}(a,b)|q_{\phi}(a,b))$
$=\mathrm{E}_{(a,b)\sim p_{\phi}(a,b)}\Big{[}D_{\mathrm{KL}}(p(i,j)|q_{\phi}(i,j|a,b))\Big{]}+D_{\mathrm{KL}}(p_{\phi}(a,b)|q_{\phi}(a,b))$
$=\mathrm{E}_{a,b\sim p_{\phi}(b|a)\mathrm{U}_{\mathcal{K}}(a)}\big{[}\mathrm{Loss}_{\rm node}^{a,b}\big{]}+\mathrm{Loss}_{\rm Aug}$
The derivation from the second to the third line utilizes the basic decomposition of KL divergence: $D_{\rm KL}(p(x,y)|q(x,y))=E_{x\sim p(x)}[D_{\rm KL}(p(y|x)|q(y|x))]+D_{\rm KL}(p(x)|q(x))$.
@@ Appendix B Algorithm
4:     $\mathcal{G}_{a},\mathcal{G}_{b}\leftarrow\text{The $a$-th and $b$-th artificial augmentation of $\mathcal{G}$}$
5:     $\mathcal{H}_{a},\mathcal{H}_{b}\leftarrow\text{Pick up the $a$-th and $b$-th knockdown teacher GRN from $\mathcal{H}_{\rm all}$}$
6:     $Z^{a},Z^{b}\leftarrow f_{\phi}(\mathcal{G}_{a}),~{}f_{\phi}(\mathcal{G}_{b})$triangleright Embedding target GRNs
7:     $Y^{a},Y^{b}\leftarrow f_{\phi}(\mathcal{H}_{a}),~{}f_{\phi}(\mathcal{H}_{b})$triangleright Embedding teacher GRNs
8:     $q_{\phi}(b|a)\leftarrow\text{softmax}\left(\Big{[}\frac{\langle Z^{a},Z^{*}\rangle_{F}}{\tau_{\rm a}}\Big{]}\right)[b]$triangleright Calculate augmentation-level target model
9:     $p_{\phi}(b|a)\leftarrow\text{softmax}\left(\Big{[}\frac{\langle Y^{a},Y^{*}\rangle_{F}}{\tau_{\rm a}}\Big{]}\right)[b]$triangleright Calculate augmentation-level teacher model
10:     $\text{Loss}_{\rm Aug}\leftarrow|\mathcal{K}|p_{\phi}(b|a)\big{(}\log p_{\phi}(b|a)-\log q_{\phi}(b|a)\big{)}$triangleright Importance sampling of $\text{Loss}_{\rm Aug}^{a,b}$
11:     $q_{\phi}(j|i,a,b)\leftarrow\text{softmax}\left(\Big{[}\frac{\langle z^{a}_{i},z^{b}_{*}\rangle}{\tau_{n}}\Big{]}\right)[j]$triangleright Calulate node-level target model
@@
13:     $\text{Loss}_{\rm SupGCL}\leftarrow|\mathcal{K}|p_{\phi}(b|a)\text{Loss}_{\rm node}^{a,b}+\text{Loss}_{\rm Aug}$triangleright Importance sampling of $\text{Loss}_{\rm node}^{a,b}$
@@ C.1 Details on the TCGA Dataset
In this study, we utilized the TCGA TARGET GTEx platform provided by UCSC Xena to extract patient-specific expression data and clinical labels (survival time data, disease subtype labels) for downstream tasks.
The TCGA platform contains four datasets: the gene expression dataset "dataset: gene expression RNAseq – RSEM norm_count," the dataset for cancer type attributes of individual patients "dataset: phenotype – TCGA TARGET GTEx selected phenotypes," the patient prognosis dataset "dataset: phenotype – TCGA survival data," and the patient phenotype data "dataset: phenotype - Phenotypes."
@@ C.2 Details on the LINCS Dataset
In this study, we used the Level 3 normalized gene expression data (filename on LINCS datasets: "GSE92742_Broad_LINCS_Level3_INF_mlr12k_n1319138x12328.gctx.gz") provided from the GEO dataset (GSE92742) of the LINCS L1000 project. Additionally, by referring to the concurrently provided experimental metadata ("GSE92742_Broad_LINCS_inst_info.txt.gz" indicating cell lines and treatment conditions, and "GSE92742_Broad_LINCS_pert_info.txt.gz" indicating drug and gene knockdown information), we extracted only the shRNA-mediated knockdown experiment groups. The cell lines and treatment durations were limited to samples from: MCF7 breast cancer cell line treated for 96 hours, A549 lung cancer cell line treated for 96 hours, and HT29 colon cancer cell line treated for 96 hours. The expression data was limited to 978 landmark genes.

### v4

@@ 5 Experiments
Figure 2: Overview of downstream tasks. Node-level tasks involve gene classification into Biological Process [BP.], Cellular Component [$C$.], and cancer relevance [Rel.]. Graph-level tasks include patient survival prediction [Hazard] and breast cancer subtyping [Subtype]. Mean pooling provides graph-level representations.
Table 1: Description of the downstream tasks
@@
[Subtype]: Disease subtype prediction Classification (5 groups) Accuracy
This chapter verifies the effectiveness of the proposed method using actual GRNs from cancer patients and biologically augmented GRNs based on gene knockdown experiments.
@@ 5.1 Benchmark of Gene Regulatory Networks
We evaluated the proposed method through the following procedure. First, we constructed patient-specific GRNs from cancer patient gene expression data and teacher GRNs from gene knockdown experiment data. Then, pre-training was performed on the proposed method using both the patient-specific and teacher GRNs. Subsequently, the performance of the downstream tasks, such as classification accuracy and regression performance, was evaluated using the pre-trained models and compared against comparative methods.
@@
: Directly performs classification or regression for downstream tasks.
@@
(Kipf & Welling, 2016): Reconstruction based graph representation learning method.
@@
(You et al., 2021): Graph contrastive learning using positive pairs between graphs.
@@
(Zhu et al., 2020): Node-level graph contrastive learning method.
@@
(He et al., 2024): State-of-the-art augmentation-free GCL.
Table 2: Fine-tuning results of node-level downstream tasks
@@
Table 3: Fine-tuning results of graph-level downstream tasks
@@
To evaluate the performance of SupGCL, we conducted benchmark evaluations using real-world datasets. For constructing patient-specific GRNs, we used cancer cell sample data from The Cancer Genome Atlas (TCGA). For constructing teacher GRNs, we used gene knockdown experiment data from cancer cell lines in the Library of Integrated Network-based Cellular Signatures (LINCS). The TCGA dataset (Weinstein et al., 2013) and the LINCS dataset  (Subramanian et al., 2017) are both large-scale and widely-used public platforms providing gene expression data from cancer patients and cell lines, respectively.
@@
The TCGA dataset also includes survival status and disease subtype labels associated with each gene expression profile. Additionally, each gene was annotated with multi-labels based on Gene Ontology  (Ashburner et al., 2000) — Biological Process (metabolism, signaling, cell organization; 3 classes), and Cellular Component (nucleus, mitochondria, ER, membrane; 4 classes). We also used the OncoKB  (Chakravarty et al., 2017) cancer-related gene list to assign binary relevance labels. These labels were used for downstream tasks. Details are provided in Appendix C.
@@
To estimate the network structure of each GRN from gene expression data, we used a Bayesian network structure learning algorithm based on B-spline regression (Imoto et al., 2002). For each experiment, gene expression values were used as node features, while edge features were defined as the linear sum of estimated regression coefficients and the parent node’s gene expression  (Tanaka et al., 2020). This structure estimation was performed per cancer type per dataset using the above algorithm. Further details are provided in Appendix D.
Figure 3: t-SNE visualization of pre-trained graph-level embeddings on breast cancer GRNs. Each point represents the readout feature of an individual patient’s network. NMI and ARI scores indicate quantitative clustering metrics of the embeddings.
@@ 5.2 Result 1: Evaluation by Downstream Task
Graph-level tasks (hazard prediction, subtype classification) used survival and subtype labels from TCGA. Note that subtype classification was conducted only for breast cancer. Node-level tasks (Biological Process - BP., Cellular Component - CC., and cancer relevance - Rel.) used gene-level annotations from Gene Ontology and OncoKB. Details of these downstream tasks are summarized in Table 1.
@@
For all methods including the SupGCL and conventional methods, we used the same 5-layer Graph Transformer architecture (Shi et al., 2021). Hyperparameters for pre-training were tuned with Optuna (Akiba et al., 2019), and the model was optimized using the AdamW optimizer (Loshchilov & Hutter, 2019). All training runs were performed on a single NVIDIA H100 SXM5 GPU. Additional experimental details can be found in Appendix E.
@@
For node-level tasks, many existing methods did not show much improvement over without-pretrain, whereas SupGCL consistently demonstrated strong performance.
In graph-level tasks like hazard and subtype prediction, while some existing methods showed marginal improvement over without-pretrain, SupGCL achieved significantly higher performance.
In addition, we evaluated the generalization performance across different cancer types. While the pre-training was robust even when the cancer types of the teacher and patient GRNs differed, the fine-tuning was not (see Appendix G for details).
5.3 Result 2: Effectiveness of Supervised Learning
In this section, to verify the effectiveness of the reference model $p_{\phi}$, we conducted additional experiments by varying the augmentation-level temperature parameter $\tau_{\rm a}$, which controls its degree of influence. This experiment also serves to empirically validate the theoretical relationship (Corollary 1) that as $\tau_{\rm a}$ increases, the SupGCL loss function asymptotically approaches that of unsupervised node-level GCL like GRACE. The experimental results for breast cancer in Table 4 show that SupGCL outperforms GRACE in all settings, demonstrating the effectiveness of incorporating biologically plausible information. Furthermore, as $\tau_{\rm a}$ increases and the influence of supervised information weakens, the performance of SupGCL tends to approach that of the unsupervised method GRACE, which supports Corollary 1. It should be noted that due to the different definitions of the denominator in their loss functions, this experiment and GRACE are not strictly identical.
Table 4: Ablation study of SupGCL on the augmentation temperature $\tau_{\rm a}$ for breast cancer.
$\tau_{\rm a}$ Hazard Subtype BP. $C$. Rel.
0.10 $\mathbf{0.670\pm 0.078}$ $0.837\pm 0.029$ $\mathbf{0.262\pm 0.035}$ $0.289\pm 0.049$ $0.586\pm 0.047$
0.25 (default) $0.650\pm 0.059$ $\mathbf{0.847\pm 0.036}$ $0.243\pm 0.052$ $\mathbf{0.291\pm 0.026}$ $0.600\pm 0.057$
0.50 $0.648\pm 0.053$ $0.846\pm 0.032$ $0.261\pm 0.034$ $0.284\pm 0.061$ $\mathbf{0.606\pm 0.039}$
1.00 $0.640\pm 0.056$ $0.835\pm 0.031$ $0.244\pm 0.042$ $0.280\pm 0.032$ $0.596\pm 0.048$
2.00 $0.656\pm 0.060$ $0.842\pm 0.031$ $0.237\pm 0.024$ $0.277\pm 0.058$ $0.590\pm 0.044$
Reference: GRACE $0.642\pm 0.064$ $0.841\pm 0.026$ $0.230\pm 0.051$ $0.236\pm 0.026$ $0.575\pm 0.035$
5.4 Result 3: Latent Space Analysis
We visualized the graph-level embedding spaces from the breast cancer dataset (Figure 3; additional visualizations including node-level embeddings are in Appendix F). The embeddings, colored by disease subtype, show that GAE and GraphCL fail to separate subtypes, while GRACE and SGRL exhibit moderate separation. SupGCL achieves the clearest separation, indicating a superior ability to learn subtype-specific representations. This qualitative observation was confirmed by quantitative clustering metrics. SupGCL consistently outperformed other methods on both the Normalized Mutual Information (NMI) and the Adjusted Rand Index (ARI), affirming that its embeddings more effectively capture subtype-specific structure.
@@ 6 Conclusion
A key limitation of our study is that SupGCL, when pre-trained on a specific cancer type, fails to improve performance on downstream tasks for other cancer types. As future work, we plan to expand the target cancer types and develop a large-scale, general-purpose SupGCL model that can operate across multiple cancer types.
Acknowledgments
This work was supported by JST Moonshot R&D Grant Number JPMJMS2021, JPMJMS2024, and JSPS KAKENHI Grant Number 25K00148. This paper is also based on results obtained from a project, JPNP14004, commissioned by the New Energy and Industrial Technology Development Organization (NEDO) This work used computational resources of the supercomputer Fugaku provided by RIKEN through the HPCI System Research Project (Project IDs: hp150272, ra000018) and TSUBAME4.0 supercomputer at Institute of Science Tokyo.
[…]
Appendix Contents
@@ Appendix A Detail Proof of Theorem 1
Since the loss function of contrastive learning (1) is formulated as the KL divergence between the joint distributions $p(x,y)$ and $q(x,y)$, under the assumption that $p(x)$ and $q(x)$ follow uniform distributions. Then, the following derived for $\mathrm{Loss}_{\rm I-CON}$:
@@
$=\mathbb{E}_{(a,b)\sim p_{\phi}(a,b)}\Big[D_{\mathrm{KL}}(p_{\phi}(i,j|a,b)|q_{\phi}(i,j|a,b))\Big]+D_{\mathrm{KL}}(p_{\phi}(a,b)|q_{\phi}(a,b))$
$=\mathbb{E}_{(a,b)\sim p_{\phi}(a,b)}\Big[D_{\mathrm{KL}}(p(i,j)|q_{\phi}(i,j|a,b))\Big]+D_{\mathrm{KL}}(p_{\phi}(a,b)|q_{\phi}(a,b))$
$=\mathbb{E}_{a,b\sim p_{\phi}(b|a)\mathrm{U}_{\mathcal{K}}(a)}\big[\mathrm{Loss}_{\rm node}^{a,b}\big]+\mathrm{Loss}_{\rm Aug}$
The derivation from the second to the third line utilizes the basic decomposition of KL divergence: $D_{\rm KL}(p(x,y)|q(x,y))=\mathbb{E}_{x\sim p(x)}[D_{\rm KL}(p(y|x)|q(y|x))]+D_{\rm KL}(p(x)|q(x))$.
@@ Appendix B Algorithm
[…]
6:  ${\bm{Z}}^{a},{\bm{Z}}^{b}\leftarrow f_{\phi}(\mathcal{G}_{a}),~f_{\phi}(\mathcal{G}_{b})$ $\triangleright$ Embedding target GRNs
7:  ${\bm{Y}}^{a},{\bm{Y}}^{b}\leftarrow f_{\phi}(\mathcal{H}_{a}),~f_{\phi}(\mathcal{H}_{b})$ $\triangleright$ Embedding teacher GRNs
8:  $q_{\phi}(b|a)\leftarrow\text{softmax}\left(\Big[\frac{\langle{\bm{Z}}^{a},{\bm{Z}}^{*}\rangle_{F}}{\tau_{\rm a}}\Big]\right)[b]$ $\triangleright$ Calculate augmentation-level target model
9:  $p_{\phi}(b|a)\leftarrow\text{softmax}\left(\Big[\frac{\langle{\bm{Y}}^{a},{\bm{Y}}^{*}\rangle_{F}}{\tau_{\rm a}}\Big]\right)[b]$ $\triangleright$ Calculate augmentation-level teacher model
10:  $\text{Loss}_{\rm Aug}\leftarrow|\mathcal{K}|p_{\phi}(b|a)\big(\log p_{\phi}(b|a)-\log q_{\phi}(b|a)\big)$ $\triangleright$ Importance sampling of $\text{Loss}_{\rm Aug}^{a,b}$
11:  $q_{\phi}(j|i,a,b)\leftarrow\text{softmax}\left(\Big[\frac{\langle{\bm{z}}^{a}_{i},{\bm{z}}^{b}_{*}\rangle}{\tau_{n}}\Big]\right)[j]$ $\triangleright$ Calculate node-level target model
@@
13:  $\text{Loss}_{\rm SupGCL}\leftarrow|\mathcal{K}|p_{\phi}(b|a)\text{Loss}_{\rm node}^{a,b}+\text{Loss}_{\rm Aug}$ $\triangleright$ Importance sampling of $\text{Loss}_{\rm node}^{a,b}$
@@ C.1 Details on the TCGA Dataset
In this study, for gene expression data, we used normalized count data from the TCGA TARGET GTEx study provided by UCSC Xena (UCSC Xena (2016)) for the TCGA dataset. For the LINCS dataset, we used normalized gene expression data from the LINCS L1000 GEO dataset (GSE92742) (Subramanian (2017)).
The TCGA platform contains four datasets: the gene expression dataset ”dataset: gene expression RNAseq – RSEM norm_count,” the dataset for cancer type attributes of individual patients ”dataset: phenotype – TCGA TARGET GTEx selected phenotypes,” the patient prognosis dataset ”dataset: phenotype – TCGA survival data,” and the patient phenotype data ”dataset: phenotype - Phenotypes.”
@@ C.2 Details on the LINCS Dataset
In this study, we used the Level 3 normalized gene expression data (filename on LINCS datasets: ”GSE92742_Broad_LINCS_Level3_INF_mlr12k_n1319138x12328.gctx.gz”) provided from the GEO dataset (GSE92742) of the LINCS L1000 project. Additionally, by referring to the concurrently provided experimental metadata (”GSE92742_Broad_LINCS_inst_info.txt.gz” indicating cell lines and treatment conditions, and ”GSE92742_Broad_LINCS_pert_info.txt.gz” indicating drug and gene knockdown information), we extracted only the shRNA-mediated knockdown experiment groups. The cell lines and treatment durations were limited to samples from: MCF7 breast cancer cell line treated for 96 hours, A549 lung cancer cell line treated for 96 hours, and HT29 colon cancer cell line treated for 96 hours. The expression data was limited to 978 landmark genes.